Cobalt crystallizes in a HCP unit cell. Its atomic radius is 0.1253nm. Its density is 8,900kg /m3. Using this information, estimate AVOGADRO's number.


1
Expert's answer
2021-04-19T05:07:30-0400

Density= 8900Kg/m³= 8.9g/cm

Radius= 0.1253nm= 1.253x10-8

One mole of Co= 59g

Volume occupied by one mole of Co= mass/density= 59/8.9 = 6.63cm3

Volume of one atom of Co= 4/3πr3

= 4/3 x 3.142x (1.253x10-8

= 8.24x10-24cm3

number of atoms present in one mole of Cobalt= volume occupied by one mole/volume occupied by one atom

= 6.63/8.24x1023

= 8.04x1023 atoms

We can see that this value is close to the Avogadro's number which is 6.02x1023 atoms.